Complete the table
2. Complete the following table: Compound MW equivalents mmo mL density Acetic acid, conc. 5.84 glacial) Isopentyl 2.92 alcohol Sulfuric (I8M) acid, conc. XXX 0.016 3. Calculate the theoretical yield for the synthesis of isopentyl acetate.
Answer #1

1 mole of acetic acid react with 1 mole of isopentyl alcohol to from product

here isopentyl alcohol is limiting reagent so

mole of product form will be = 2.92 mmoles

molar mass of isopentyl acetate ( product) = 130.18 g/mol

theroretical yield = 2.92 * 130.18 = 380 mg

theoretical yield = 0.38 g

table   

compound MW equivalent mmol g mL

acetic acid 60.05 g/mol 5.84/2.92 = 2 5.84 0.351

isopentyl alcohol 88.148 g/mol 2.92/2.92 = 1 2.92 0.257

sulfuric acid (18M) 0.288/2.92=0.1 0.288 0.016

amount of equivalent is calculated with respect to the stoichiometric amount of limiting reagent.